### v2

@@ 4.1 Second-order uncertainty quantification methods
Bayesian methods form a principled approach to quantify and interpret epistemic uncertainty by modelling a prior and posterior distribution over the model parameters and computing the second-order distribution via Eqn. 1. For the regression setup that we consider in this section, a common approach is to use Heteroscedastic Gaussian Processes (GP, [19, 41]), which place priors on the parameters $f(\boldsymbol{x})$ and $\sigma^{2}(\boldsymbol{x})$, thereby analytically computing the predictive distribution using two Gaussian processes. For Bayesian neural networks, which have gained popularity in recent years, the posterior and second-order distribution cannot be analytically computed, due to the intractability of $p(\boldsymbol{\phi}|\mathcal{D}_{N})\propto p(\mathcal{D}_{N}|\boldsymbol{\phi})p(\boldsymbol{\phi})$ induced by the hypothesis space [1]. Due to the impossibility of performing exact inference in Bayesian neural networks, one typically obtains $d$ different samples $\{\boldsymbol{\phi}_{j}\}_{j=1}^{d}$ from $p(\boldsymbol{\phi}|\mathcal{D}_{N})$ by use of approximation techniques that aim to find a posterior distribution via sampling methods. Here, the second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$ of Eqn. 1 is then approximated by Monte Carlo integration:
$q(\boldsymbol{\theta}|\boldsymbol{x})=\mathbb{E}_{\boldsymbol{\phi}\sim p(\boldsymbol{\phi}|\boldsymbol{x})}\big[q(\boldsymbol{\theta}\mid\boldsymbol{x},\boldsymbol{\phi})\big]\approx\frac{1}{d}\sum_{j=1}^{d}q(\boldsymbol{\theta}\mid\boldsymbol{x},\boldsymbol{\phi}_{j}).$ (9)
In the experiments below, we employ three of the most common methods to obtain $\{\boldsymbol{\phi}_{i}\}_{j=1}^{M}$. Variational inference, (VI, [42, 43]) considers a variational (simpler) distribution $q(\boldsymbol{\phi})$ and optimizes the lower bound of the marginal likelihood (ELBO). Hamiltonian Monte Carlo (HMC) uses Hamiltonian dynamics to sample $\{\boldsymbol{\phi}_{j}\}$ from the true posterior. Laplace approximation (LA, [44]) builds a Taylor expansion around the maximum-a-posteriori estimate $\boldsymbol{\phi}_{\text{MAP}}$, yielding a Gaussian approximation. Ensemble methods are also often interpreted as performing approximate Bayesian inference by finding multiple modes of $p(\boldsymbol{\phi}|\mathcal{D}_{N})$ after training multiple independent models. [45, 46]. In this work, we consider three types of ensembles. Deep Ensembles (DE, [9]), train $d$ networks on the same dataset $\mathcal{D}_{N}$, but with different random initialization of the weights $\boldsymbol{\phi}$. Similarly, in Bootstrap Ensembles (BE), $M$ networks are trained using different weight initializations, but each network uses a bootstrap sample $\mathcal{D}^{(j)}\subseteq\mathcal{D}_{N}$. While Bootstrap Ensembles have been shown to perform worse than training $d$ networks using Deep Ensembles [9], we include them in our analysis to estimate the effect of data uncertainty on the epistemic uncertainty estimates. Monte Carlo Dropout (MCD, [47]) can also be understood as a Bayesian approximation, where one trains a single neural network with dropout rate $\rho$ and performs multiple stochastic forward passes at inference time with the same dropout rate. In a different direction, evidential deep learning methods learn directly the second-order distribution over the predictive parameters $\theta$. Deep evidential regression (DER, [12]) learns hyperparameters of a conjugate prior (e.g. Normal-Inverse-Gamma for the univariate normal distribution) directly using a single network trained by second-order loss minimization.
4.2 Uncertainty disentanglement
From a second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$, one can decompose its representation of uncertainty into an estimate of aleatoric and epistemic uncertainty. Proposed measures focus on entropy [3], distance to reference sets [10], or proper scoring rules [48, 49]. The most commonly used one for regression is arguably the variance-based decomposition [3, 10], which considers the variance of the prediction $\hat{y}$, obtained via model averaging as in Eqn. 1. The latter can be decomposed into (estimates of) an aleatoric and an epistemic part:
$\text{Var}(\hat{y}|\boldsymbol{x})=\underbrace{\mathbb{E}_{\boldsymbol{\theta}\sim q(\boldsymbol{\theta}\mid\boldsymbol{x})}\big[\text{Var}(\hat{y}|\boldsymbol{x},\boldsymbol{\theta})\big]}_{\text{ aleatoric estimate}}+\underbrace{\text{Var}_{\boldsymbol{\theta}\sim q(\boldsymbol{\theta}\mid\boldsymbol{x})}\big(\mathbb{E}[\hat{y}|\boldsymbol{x},\boldsymbol{\theta}]\big)}_{\text{epistemic estimate}}$ (10)
Here, the first term estimating aleatoric uncertainty describes the average predicted variance $\sigma^{2}$. The second term captures the variance of the mean prediction $f$ and hence an estimate of epistemic uncertainty. Note here that, unlike in the bias-variance decomposition in Eqn. 2.2, the bias in the prediction is not represented by this decomposition. Since any second-order distribution does not incorporate ground-truth knowledge, any systematic bias goes unaccounted. Moreover, second-order uncertainty quantification methods lack informativeness about which source of epistemic uncertainty they capture. In the next paragraph we illustrate these two concerns for two methods that serve as representative examples.
4.3 Two representative cases
Example 1 (Deep Ensembles purely capture procedural uncertainty.).
Let $\gamma\sim P^{\gamma}$ denote the random seed (weight initialization, data ordering etc.) that differentiates each ensemble member’s training, which in turn are trained on the same dataset $\mathcal{D}_{N}$. For a Deep Ensemble of size $d$, each ensemble member’s predictive parameters are indexed over the set of random seeds $\{\gamma_{1},\dots,\gamma_{d}\}$, obtaining a set of predictions $\{\boldsymbol{\theta}_{\gamma_{i}}(\boldsymbol{x})=(f_{\gamma_{i}}(\boldsymbol{x}),\sigma_{\gamma_{i}}^{2}(\boldsymbol{x}))\}_{i=1}^{d}$. The empirical second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$ can be approximated as a mixture of equally weighted Dirac measures
$q(\boldsymbol{\theta}|\boldsymbol{x})=\frac{1}{d}\sum_{i=1}^{d}\delta_{\boldsymbol{\theta}_{\gamma_{i}}(\boldsymbol{x})}(\boldsymbol{\theta}),$
[…]
$\text{Var}(\hat{y}|\boldsymbol{x})$ $=\frac{1}{d}\sum_{i=1}^{d}\sigma_{\gamma_{i}}^{2}(\boldsymbol{x})+\frac{1}{d}\sum_{i=1}^{d}(f_{\gamma_{i}}(\boldsymbol{x})-\hat{f}(\boldsymbol{x}))^{2}$
$\longrightarrow$ $\,\mathbb{E}_{\gamma}\big[\text{Var}(\hat{y}|\boldsymbol{x},\gamma)\big]+\text{Var}_{\gamma}\big(\mathbb{E}_{y|\boldsymbol{x}}[\hat{y}|\boldsymbol{x},\gamma]\big),\,d\rightarrow\infty.$
Here, the first part corresponds to the aleatoric uncertainty estimate and the epistemic uncertainty estimate is given by
$\text{Var}_{\gamma}\big(\mathbb{E}_{y|\boldsymbol{x}}[\hat{y}|\boldsymbol{x},\gamma]\big)=\mathbb{E}_{\gamma}\big[(\mathbb{E}_{x|y}[\hat{y}|\boldsymbol{x},\gamma]-\mathbb{E}_{\gamma}[\mathbb{E}_{y|\boldsymbol{x}}[\hat{y}|\boldsymbol{x},\gamma]])^{2}\big]=\mathbb{E}_{\gamma}\big[(\hat{y}_{N,\gamma}(\boldsymbol{x})-\hat{y}_{N}(\boldsymbol{x}))^{2}\big],$
where $\hat{y}_{N,\gamma}(\boldsymbol{x}):=f_{\gamma}(\boldsymbol{x})$ corresponds to the prediction of an individual ensemble member trained with random initialization $\gamma$ on $\mathcal{D}_{N}$ and $\hat{y}_{N}(\boldsymbol{x}):=\hat{f}(\boldsymbol{x})$ of the aggregated ensembles. Hence, we see that the predicted epistemic uncertainty corresponds to the procedural uncertainty as illustrated in Figure 1(b). Model uncertainty through bias as well as data uncertainty are not represented.
While Deep Ensembles might explore multiple (local) optima of the training loss by using different random seeds [45], Bayesian approaches attempt to explicitly approximate the posterior over parameters, which can yield more principled estimates of epistemic uncertainty. However, the practical methods used to approximate the posterior distribution also face restrictions: Approximate inference methods like the Laplace approximation or variational inference often capture only a single local mode [45], employ simplifying assumptions [25], or can be computationally costly in high-dimensional weight spaces. Other issues arise with methods that try to estimate the second-order distribution directly via loss minimisation, such as Deep Evidential Regression.
Example 2 (DER may fail to provide faithful epistemic uncertainty estimates.).
Deep Evidential Regression [12] learns to predict the parameters of a Normal–Inverse-Gamma second–order distribution
$q_{\text{NIG}}(\boldsymbol{\theta}|\boldsymbol{x})=\mathrm{NIG}\!\bigl(\gamma(\boldsymbol{x}),\nu(\boldsymbol{x}),\alpha(\boldsymbol{x}),\beta(\boldsymbol{x})\bigr),$
which forms the conjugate prior of the normal distribution. The parameters $(\gamma,\nu,\beta,\alpha)$ are learned by minimising a second-order loss, i.e. the loss function one derives when doing Bayesian inference with the NIG prior, together with a regularization term called the evidence, which controls the uncertainty for data points far from the training set. Using the variance-based decomposition, one obtains analytical expressions for the aleatoric and epistemic uncertainty estimates, namely
$\mathbb{E}_{\boldsymbol{\theta}\sim q_{\text{NIG}}(\boldsymbol{\theta}|\boldsymbol{x})}[\text{Var}(\hat{y}|\boldsymbol{x},\boldsymbol{\theta})\big]=\frac{\beta}{\alpha-1},\quad\text{Var}_{\boldsymbol{\theta}\sim q_{\text{NIG}}(\boldsymbol{\theta}|\boldsymbol{x})}\big(\mathbb{E}_{y|\boldsymbol{x}}[\hat{y}|\boldsymbol{x},\boldsymbol{\theta}]\big)=\frac{\beta(\alpha-1)}{\nu}.$
However, a line of work has shown that the method has severe drawbacks in the optimization [27, 50], as well as the estimation of aleatoric and epistemic uncertainty [51], and the representation of the second-order distribution [24], making it more of a heuristic than a principled method.
4.4 Synthetic data illustration

### v3

@@ 4.1 Second-order uncertainty quantification methods
The prevailing approach represents epistemic uncertainty through second-order distributions over predictive distributions, from which aleatoric and epistemic components are extracted via decomposition formulas (Hüllermeier and Waegeman, 2021; Gawlikowski et al., 2023). Bayesian neural networks (Neal, 2012) form a principled approach to quantify epistemic uncertainty by modelling a prior and posterior distribution over the neural network’s weights. Due to the intractability of the posterior distribution induced by this hypothesis space (Kendall and Gal, 2017), one typically samples from it, using different approximation techniques. The second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$ is then approximated via Monte Carlo integration. Variational inference (Jordan et al., 1999; Blundell et al., 2015; Blei et al., 2017) considers a variational (simpler) distribution and optimizes the lower bound of the marginal likelihood. However, posterior quality depends critically on the variational approximation and prior specification. Monte Carlo Dropout (Gal and Ghahramani, 2016) can also be understood as a Bayesian approximation, where one trains a single neural network with dropout rate $\rho$ and performs multiple stochastic forward passes at inference time with the same dropout rate. Spectral-Normalized Gaussian Processes (Liu et al., 2020a) represent the second-order distribution by approximating a Gaussian distribution over the classifier output. Stochastic Weight Averaging Gaussian (Maddox et al., 2019) fits a Gaussian distribution over the model weights by using the stochastic gradient descent iterates to derive estimates for moment and covariance. Similarly, the Laplace approximation (MacKay, 1992; Daxberger et al., 2021) yields a Gaussian approximation by building a Taylor expansion around the maximum-a-posteriori estimate. Deep Ensembles (Lakshminarayanan et al., 2017) train a number of different models on the same dataset, but with different weight initializations. The resulting predictions can be seen as samples from an approximate posterior, thereby forming a Bayesian approximation. In Evidential Deep Learning (Ulmer et al., ; Sensoy et al., 2018; Malinin and Gales, 2018b), the parameters of the second-order distribution are directly learned via loss minimization. Sensoy et al. (2018) place Dirichlet distributions over class probabilities for classification, while Deep Evidential Regression (Amini et al., 2020) extends this to regression using the Normal-Inverse-Gamma as conjugate prior distribution.
4.2 Alternative uncertainty representations
Given the documented limitations of second-order distributions, several alternative frameworks have emerged. We focus on deterministic methods based on distances or direct loss prediction, which have gained traction for computational efficiency and competitive empirical performance, then briefly discuss credal sets and conformal prediction.
Deterministic methods. Deterministic uncertainty quantification methods (Postels et al., 2022) quantify epistemic uncertainty using either the distribution of latent representations learned by the model, or by directly predicting the loss of the network’s predictions. Deterministic Uncertainty Quantification (Van Amersfoort et al., 2020) replaces the softmax layer with an RBF-like output layer around class centroids, thereby measuring distance to the centroids as a proxy for uncertainty. The Mahalanobis method (Lee et al., 2018) captures epistemic uncertainty post-hoc by measuring the distance from the learned class-conditional distributions in feature space. Direct Epistemic Uncertainty Prediction (Lahlou et al., 2023) quantifies epistemic uncertainty as the excess risk of the predictor, i.e. the gap between pointwise expected loss and the loss of the Bayes predictor. It is learned via a secondary model trained on out-of-sample errors. Distance-based methods implicitly capture a form of distributional uncertainty: distance from training data correlates with regions where predictions are unreliable. However, they do not explicitly decompose epistemic uncertainty into data and procedural components, nor do they account for systematic bias. On the other hand, methods based on predicting the loss of the model are robust to model bias, but require training and maintaining an additional error-prediction model.
Credal sets and imprecise probabilities. Rather than placing distributions over distributions, credal sets represent epistemic uncertainty as convex sets of probability distributions (Walley, 1991). Recent work has operationalized credal sets in deep learning (Caprio et al., 2024b, a). Theoretically, credal sets offer a more expressive uncertainty language, however, they are not yet fully established in mainstream machine learning practice. Computational challenges arise from working with sets rather than point estimates, and the connection to standard training procedures remains less direct than for probabilistic methods.
Conformal prediction. Conformal prediction (Vovk et al., 2005; Angelopoulos et al., 2023) produces prediction sets with finite-sample coverage guarantees, which hold for any sample size and any underlying model. However, it addresses a different problem than epistemic uncertainty quantification, since it does not decompose uncertainty into aleatoric and epistemic components. Conformal methods are thus complementary to, rather than a replacement for, principled epistemic uncertainty estimation.
5 Conclusion and call to action
Epistemic uncertainty is generally not well defined and is still a heuristic concept for machine learning models, leading to interpretability issues for uncertainty estimates. Our analysis reveals that the prominent second-order uncertainty estimation methods systematically fail to capture what they claim to measure. Two mechanisms drive this failure: First, the predictive bias contaminates the aleatoric-epistemic uncertainty decomposition and should be evaluated as an epistemic component. Second, epistemic uncertainty estimates typically only reflect parts of the true variance, often capturing procedural rather than data uncertainty.
Call to action. We advocate for a more fine-grained taxonomy that explicitly distinguishes between model uncertainty, comprising approximation error plus estimation bias, and estimation uncertainty, comprising data and procedural components. New methods should be evaluated not only on downstream tasks but also against reference distributions that decompose variance into different sources. Claims of epistemic uncertainty estimation should specify which components are captured and acknowledge which are not. In safety-critical applications, we recommend: (1) assessing model bias separately through held-out validation or domain expertise, (2) treating epistemic uncertainty estimates as lower bounds on reducible uncertainty rather than complete characterizations, and (3) using multiple complementary uncertainty methods, as different methods capture different uncertainty components.
Appendix A Impact of recent literature
In recent years, uncertainty disentanglement has gained importance in the machine learning literature. Considering the following collection of highly relevant papers, they have received almost 40,000 citations according to Google Scholar: (Der Kiureghian and Ditlevsen, 2009; Kendall and Gal, 2017; Hüllermeier and Waegeman, 2021; MacKay, 1992; Depeweg et al., ; Malinin and Gales, 2018a, 2019; Amini et al., 2020; Lakshminarayanan et al., 2017; Wilson and Izmailov, 2020; Gal and others, 2016; Charpentier et al., 2020, 2021; Stadler et al., 2021; Lee et al., 2018; Liu et al., 2020b; Van Amersfoort et al., 2020; Sun et al., 2022; Deng et al., 2023)
Appendix B Incomplete uncertainty estimators
B.1 Bias aleatoric uncertainty estimators
For a data-generating process such that $y_{i}=f(\boldsymbol{x}_{i})+\epsilon_{i}$, where $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}(\boldsymbol{x}_{i}))$ is data-dependent (heteroscedastic) noise, and $f$ is a deterministic function, such that $p(\cdot|\boldsymbol{x})$ forms the density of a Gaussian distribution with mean $f(\boldsymbol{x})$ and variance $\sigma^{2}(\boldsymbol{x})$. There are multiple ways to model aleatoric uncertainty. One popular way in the literature to do so (Zhang et al., 2024) is using the variance attenuation (Lakshminarayanan et al., 2017; Amini et al., 2020; Valdenegro-Toro and Mori, 2022), which means to train a neural network with two heads (one that outputs the predictive mean and the other that outputs the predictive variance) using a loss function that reflects the likelihood of the observed data:
$\ell_{va}=\frac{1}{M}\sum^{N}_{i=1}\frac{(y_{i}-\hat{f}(x_{i}))^{2}}{2\hat{\sigma}^{2}(x_{i})}+\frac{\log(\hat{\sigma}^{2}(x_{i}))}{2}.$ (9)
We rely on the derivation presented by (Zhang et al., 2024) to show how Eqn. 9 does not produce an unbiased estimator of the aleatoric uncertainty $\sigma^{2}(\boldsymbol{x})$, and refer the reader to its paper for all the details. In this setting, the expected value of the predictive variance becomes
$\mathbb{E}[\hat{\sigma}^{2}(x)]=(f(x)-\hat{f}(x))^{2}+\sigma^{2}(x).$ (10)
Where one can see that the expected aleatoric noise differs from the true aleatoric noise by a square difference of the predictive mean and the true label.
While the original contribution of (Nix and Weigend, 1994) suggest to perform a warm-up period, one can see that this modification to the training process would improve the mean prediction, by fitting the model in a simpler loss landscape, and therefore the accuracy of $\hat{f}$, leaving the expected value of the predictive variance unchanged. Other methods have been proposed to produce a better estimate of aleatoric uncertainty. (Seitzer et al., 2022) proposed adding an extra term $\lfloor\hat{\sigma}^{2\beta}(x)\rfloor$ to the second term of $\ell_{va}$, where $\lfloor\cdot\rfloor$ denotes the stop gradient operation. However, this modification does not change the expected value of the estimator, it just restricts the learning path of the predictive mean. (Sluijterman et al., 2025) proposed a method that divides the training process into two parts, by creating some modified datapoints for the training of the second part. This process improves mean predictions and the potential to avoid overfitting, while the bias of the expected aleatoric uncertainty estimation is still there when the predictive mean deviates from the true value.
B.2 Regularization effects on predictive variance
Neural networks are trained with multiple sources of regularization due to the limited sample size (e.g. early stopping, weight decay, dropout). A body of works in the literature, from textbooks to recent machine learning papers (Larsen and Hansen, 1994; Nusrat and Jang, 2018; Ninness and Hjalmarsson, 2004; Girosi et al., 1995), show how regularization reduces predictive variance by constraining how strongly the learned model can respond to finite-sample fluctuations in the training data, which happens through parameter shrinkage, effective dimensionality reduction, and smoothing of the learned function.
By reducing the variance, the third term of Eqn. 3.1 shrinks and the second term of the same equation increases, as bias increases from regularization. Our argument is key in this aspect: if we aim to capture epistemic uncertainty, looking only at the variance, which has been affected by the training process, gives an incomplete picture of the whole epistemic uncertainty.
Appendix C Experimental details
C.1 Reference distribution approximation
Synthetic dataset.
To obtain an empirical approximation of the reference distribution (see Definition  3.1), we sample $20$ different datasets from the known data-generating process to approximate the data uncertainty component of the epistemic uncertainty. For each dataset, we train $10$ different MLP models with different random weight initializations, which allows us to obtain a procedural uncertainty estimate. Each MLP uses the whole dataset. The MLP models have $4$ hidden layers with $100$ neurons each. We use a learning rate of 0.009 and train the models for $500$ epochs. We repeated the process for different training dataset sizes: $50,100,500$ with batch sizes $32,32,64$ respectively. Seeds for the datasets are $\{7,42,2024,123,999,50,100,150,200,250,300,350,400,450,500,550,600,650,700,750\}$.
Figure 5: Estimated procedural and data uncertainty of the reference distribution for increasing sample sizes. The reference distribution is estimated by resampling training data and procedural parameters, using $n_{\gamma}=10$ and $n_{d}=20$. Procedural and data uncertainty are calculated using the decomposition formula in Eqn. 3.1. Even when the model predictions (complete line) have an important deviation (bias) from the ground-truth expected value (dashed line), the epistemic uncertainty estimate decreases with more data.
Figure 5 shows the obtained procedural and data uncertainty of the reference distribution for different training dataset sizes. First, we see that the total variance of the reference distribution decreases when sample size increases, supporting the claim that the reference distribution properly represents data uncertainty. Second, we see that the estimate of the procedural uncertainty also decreases with an increasing sample size, which is not unexpected, because initial procedural configurations become less important when more data is observed. Third, in the region with high epistemic uncertainty, the reference distribution has a very large bias, due to the underlying complexity of the data-generating function. The bias is in fact, so high that aleatoric uncertainty cannot be correctly estimated anymore. This last finding confirms our central position: Any claim to properly represent epistemic uncertainty must examine not only how different sources of uncertainty are modeled, but also whether bias is systematically distorting the partitioning of uncertainty. When the bias is high, estimates of both epistemic and aleatoric uncertainty can be completely misleading.
Real dataset.
To estimate the reference distribution as described in Section 3.2, we partition the training dataset into $100$ disjoint subsets of $12,067$ observations each (i.e., $n_{d}=100$). This partition allows us to approximate the data uncertainty component from the epistemic uncertainty. We argue that in big datasets, where enough disjoint subsamples can be taken, it is possible to approximate the variance coming from different datasets at training time. For each subset, we train $n_{\gamma}=5$ multilayer perceptrons (MLPs) with different random weight initializations, this part of the process is the same as in the synthetic data setting and results in an approximation of the procedural component of the epistemic uncertainty. This whole process yields a total of $500$ independently trained models. Each model produces predictive means $\hat{y}_{N,\gamma}(\boldsymbol{x})$, and by computing the empirical variance across these 500 predictions, we obtain an estimate of $\text{Var}_{\mathcal{D}_{N},\gamma}(\hat{y}|\boldsymbol{x})$ as defined in Eqn. 8. This procedure operationalizes the reference distribution introduced in Definition 3.1: the pseudo-resampling across datasets ($\mathcal{D}_{N}$) captures data uncertainty while the random initializations account for procedural uncertainty. Given the large sample size of the dataset, the partitioning approach provides a computationally feasible approximation of repeated sampling from the underlying data-generating process, effectively emulating access to its stochastic variability. This design allows us to decompose the epistemic uncertainty component into its data and procedural parts in a controlled manner, while being in a real-world data setup.
C.2 Synthetic data experiments
To illustrate the limitations of existing uncertainty quantification methods in producing reliable disentangled estimates, we showcase the behavior of eight representative methods of the three second-order quantification approaches we discuss in the main paper. This section shows the experimental details and results from the uncertainty quantification methods applied to the problem described by Eqn. 6.
C.2.1 Experimental illustrations
We use the variance-based split of the predictive variance as in Eqn. 8 to evaluate the methods in the heteroscedastic regression setting. Figure 6 illustrates that epistemic uncertainty estimates decrease monotonically with sample size, consistent with the behavior observed for the reference distribution. This trend confirms that, as the dataset grows, models become increasingly confident, yet not necessarily more accurate, particularly in difficult-to-learn regions. In these settings, even with abundant data, the model can remain overconfident in systematically biased predictions. Such behavior reveals a key limitation of current epistemic uncertainty measures and disentanglement frameworks: they fail to account for systematic bias, leading to an underestimation of epistemic uncertainty. This pattern is not unique to Deep Ensembles, but is consistently observed across other methods (see Figures 1117), indicating that this limitation is intrinsic to the way epistemic uncertainty is commonly represented rather than a method-specific artifact. Bias is approximated as specified in Eqn. 3.1: datasets are repeatedly resampled from the true data‐generating process, models are trained on each sample, and the mean deviation of the predictions from the true conditional mean is computed. Table 1 shows that bias is frequently larger than –and largely uncorrelated with– the corresponding epistemic uncertainty estimates. These findings underscore our central position: any attempt to disentangle aleatoric from epistemic uncertainty must assess not only how each component is modeled, but also the extent to which systematic bias distorts their partitioning.